# Lie algebra homology with coefficients tensor products of the adjoint representation in relative polynomial degree $2$

### v1

@@ 4.1. The functors $\Lambda^{3}\otimes\Lambda^{1}$ and $\Lambda^{2}\otimes\Lambda^{2}$
To conclude, it suffices to observe that $\Lambda^{2}\otimes\Lambda^{2}$ is not invariant under the $\mathfrak{S}_{2}$-action transposing factors. This implies that the remaining factor must correspond to $\mathbf{S}_{(2,1^{2})}\boxtimes\mathrm{sgn}_{2}$. (This can also be verified by an explicit calculation.) ∎
@@
We record the following application of Schur’s lemma to endomorphisms of $\Lambda^{3}\otimes\Lambda^{1}$ analogous to 10 above:
@@ 4.2. The morphisms $\kappa$ and $\tilde{\kappa}$
The definition of $\tilde{\kappa}$ in terms of $\kappa$ ensures that $(\kappa,\tilde{\kappa})$ is $\mathfrak{S}_{2}$-equivariant. The result then follows from the characterization of $\overline{\delta}_{2}$. ∎
More generally, as for the description of $\delta_{r}^{\prime\prime}$ given above, one has:
Lemma 4.6.
given by $\kappa\otimes\mathrm{Id}_{T^{r-2}}$.
@@ 4.3. The isomorphism $\alpha$
Before introducing $\alpha$, we require to establish some properties of $\kappa$ and $\tilde{\kappa}$:
@@
$\kappa$ ::: $\Lambda^{3}\otimes\Lambda^{1}\rightarrow\Lambda^{2}\otimes\Lambda^{2}$
$\tilde{\kappa}$ ::: $\Lambda^{1}\otimes\Lambda^{3}\rightarrow\Lambda^{2}\otimes\Lambda^{2}$
are injective. Moreover $\mathrm{image}(\kappa)=\mathrm{image}(\tilde{\kappa})$.
@@
Under the isomorphism (11), the automorphism $\tau\circ\alpha\in\mathrm{End}(\Lambda^{3}\otimes\Lambda^{1})$ identifies as $(1,-1)$. In particular, $\tau\circ\alpha$ is an involution, i.e., $(\tau\circ\alpha)\circ(\tau\circ\alpha)=\mathrm{Id}$.
@@ 4.4. Extending $\alpha$ for $r>3$
Example 4.12.
For $r=2$, one has $\alpha_{1;2}=\alpha$. If $r>2$, then $\alpha_{1:2}$ is $\alpha\otimes\mathrm{Id}_{T^{r-2}}$.
@@
    By Proposition 4.9, we have $\alpha_{j;i}^{-1}=\tau_{i;j}\circ\alpha_{j;i}\circ\tau_{i:j}$, where $\tau_{i;j}$ transposes the $i$th and $j$th tensor factors.
@@ 4.5. Identifying $\ker\overline{\delta}_{r}$
Suppose that $r\geq 2$. Motivated by Proposition 3.9, we seek to calculate $\ker\overline{\delta}_{r}$, where $\overline{\delta}_{r}$ is the morphism introduced in Notation 3.7. The latter has the form given in Lemma 3.8:
@@
Notation 4.14.
@@
By Lemma 4.6, the definition of the $\alpha_{i,j}$ ensures that this result holds. Namely, by definition of $\overline{\delta}_{r}$, $(X_{i})_{1\leq i\leq r}$ lies in the kernel of $\overline{\delta}_{r}$ if and only if, for each $1\leq i<j\leq r$, the map
$(T^{i-1}\otimes\Lambda^{3}\otimes T^{r-i})\ \oplus\ (T^{j-1}\otimes\Lambda^{3}\otimes T^{r-j})\rightarrow T^{i-1}\otimes\Lambda^{2}\otimes T^{j-i-1}\otimes\Lambda^{2}\otimes T^{r-j}$
induced by $\kappa$ and $\tilde{\kappa}$ applied with respect to the $i$th and $j$th tensor factors is zero. Here $X_{i}$ lies in $T^{i-1}\otimes\Lambda^{3}\otimes T^{r-i}$ and $X_{j}$ in $T^{j-1}\otimes\Lambda^{3}\otimes T^{r-j}$. The latter condition is equivalent to $\alpha_{i;j}(X_{i})=-X_{j}$, by definition of $\alpha_{i;j}$. ∎
@@
    the underlying inclusion $\mathbf{S}_{(2,1^{2})}\hookrightarrow\Lambda^{3}\otimes\Lambda^{1}\oplus\Lambda^{1}\otimes\Lambda^{3}$ corresponds to a commutative diagram:
@@
By the discussion preceding the statement and Pieri’s rule, we already know that the underlying functor of $\ker\overline{\delta}_{2}$ is $\Lambda^{4}\oplus\mathbf{S}_{(2,1^{2})}$. It remains to consider the $\mathfrak{S}_{2}$-action. This is analysed by using Proposition 4.9, which shows that $\tau\circ\alpha$ corresponds to the action of the diagonal matrix with entries $(1,-1)$. The explicit description of the underlying inclusions follows directly from this. ∎
Remark 4.19.
The sign in the equality $\alpha(X_{1})=-X_{2}$ appears in different ways for the cases $\Lambda^{4}$ and $\mathbf{S}_{(2,1^{2})}$. For $\Lambda^{4}$, it arises from the ‘graded commutativity’ of the coproduct of $\Lambda^{*}$. For $\mathbf{S}_{(2,1^{2})}$, the sign arises from $\tau\circ\alpha$ whereas the embedding of $\mathbf{S}_{(2,1^{2})}$ in $\Lambda^{1}\otimes\Lambda^{3}$ is given by the commutative triangle in the statement of the Theorem, using $\tau$.
@@ 4.6. A lower bound for $\ker\overline{\delta}_{r}$
The terms $\Lambda^{r+2}\boxtimes\mathrm{sgn}_{r}$ and $\mathbf{S}_{(r,1^{2})}\boxtimes\mathrm{triv}_{r}$ are simple and non-isomorphic, hence it suffices to exhibit the appropriate inclusions for the individual factors.
For the underlying functor $\Lambda^{r+2}$, for each $i\in\{1,\ldots,r\}$, the natural transformation
$\Lambda^{r+2}\rightarrow T^{i-1}\otimes\Lambda^{3}\otimes T^{r-i}$
is given by the coproduct of $\Lambda^{*}$. Generalizing the case $r=2$ given in Theorem 4.18, one checks using coassociativity of the coproduct of $\Lambda^{*}$ that these induce an inclusion $\Lambda^{r+2}\subset\ker\overline{\delta}_{r}$. Moreover, taking into account the $\mathfrak{S}_{r}$-action, one checks that this defines a $\mathfrak{S}_{r}$-equivariant inclusion $\Lambda^{r+2}\boxtimes\mathrm{sgn}_{r}\subset\ker\overline{\delta}_{r}$.
Similarly, for $\mathbf{S}_{(r,1^{2})}$, one starts from a chosen inclusion (unique up to non-zero scalar by Schur’s Lemma) $\mathbf{S}_{(r,1^{2})}\hookrightarrow\Lambda^{3}\otimes\Gamma^{r-1}$. Observe that, by Example 2.9 and the characterization of $\mathsf{d}$ given in Lemma 2.7, this fits into a commutative diagram
where the right hand vertical map is induced by $\mathbf{S}_{(2,1^{2})}\subset\Lambda^{3}\otimes\Gamma^{1}$ and the lower horizontal map by the coproduct of $\Gamma^{*}$.
For $1\leq i\leq r$, using the coproduct of $\Gamma^{*}$, one has the composite inclusion:
$\mathbf{S}_{(r,1^{2})}\hookrightarrow\Lambda^{3}\otimes\Gamma^{r-1}\rightarrow\Lambda^{3}\otimes\Gamma^{i-1}\otimes\Gamma^{r-i}\cong\Gamma^{i-1}\otimes\Lambda^{3}\otimes\Gamma^{r-i}\subset T^{i-1}\otimes\Lambda^{3}\otimes T^{r-i}$
where the isomorphism is given by the symmetric monoidal structure and the inclusion is given by $\Gamma^{*}\subset T^{*}$. Using the commutative diagram (23) and the coassociativity of the coproduct of $\Gamma^{*}$, one deduces from Theorem 4.18 that these maps define an inclusion $\mathbf{S}_{(r,1^{2})}\subset\ker\overline{\delta}_{r}$. Moreover, from the construction, it is clear that this corresponds to a $\mathfrak{S}_{r}$-equivariant inclusion $\mathbf{S}_{(r,1^{2})}\boxtimes\mathrm{triv}_{r}\subset\ker\overline{\delta}_{r}$. ∎
5. The case $r=3$
A key step in proving Theorem 1 is the case $r=3$:
Theorem 5.1.
There is a $\mathfrak{S}_{3}$-equivariant natural isomorphism
The proof uses Proposition 4.15 for the case $r=3$. The relevant morphisms (see Notation 4.11) fit into the (non-commutative) diagram:
The cocycle condition arising from Proposition 4.15 can be restated as
Proposition 5.2.
The kernel of $\overline{\delta}_{3}$ is isomorphic to the equalizer of
The composite $\alpha_{1;3}^{-1}\circ\alpha_{2;3}\circ\alpha_{1;2}$ is an automorphism of $\Lambda^{3}\otimes\Lambda^{1}\otimes\Lambda^{1}$. This will be analysed using:
Notation 5.3.

### v4

@@ 4.1. The functors $\Lambda^{3}\otimes\Lambda^{1}$ and $\Lambda^{2}\otimes\Lambda^{2}$
To conclude, it suffices to observe that $\Lambda^{2}\otimes\Lambda^{2}$ is not invariant under the $\mathfrak{S}_{2}$-action transposing factors. This implies that the remaining factor must correspond to $\mathbf{S}_{(2,1^{2})}\boxtimes\mathrm{sgn}_{2}$. ∎
@@
We record the following application of Schur’s lemma to endomorphisms of $\Lambda^{3}\otimes\Lambda^{1}$ analogous to (12) above:
@@ 4.2. The morphisms $\kappa$ and $\tilde{\kappa}$
The significance of $\kappa$ is shown by the following Lemma, which is an addendum to Lemma 3.9:
Lemma 4.5.
given by $\kappa\otimes\mathrm{Id}_{T^{r-2}}$.
Remark 4.6.
There is a slight subtlety here: $\kappa$ is defined using the coproduct $\Lambda^{3}\rightarrow\Lambda^{2}\otimes\Lambda^{1}$, whereas $\overline{\delta}_{2}$ is defined using the coproduct $\Lambda^{3}\rightarrow\Lambda^{1}\otimes\Lambda^{2}$. This makes no difference, since the ‘graded commutativity’ of the exterior algebra in this case ensures that the two maps identify via the transposition of the tensor factors of $\Lambda^{1}\otimes\Lambda^{2}$.
@@
[…]
@@ 4.3. The isomorphism $\alpha$
Before introducing the announced isomorphism $\alpha$, we require to establish some properties of $\kappa$ and $\tilde{\kappa}$:
@@
$\kappa$ $:$ $\Lambda^{3}\otimes\Lambda^{1}\rightarrow\Lambda^{2}\otimes\Lambda^{2}$
$\tilde{\kappa}$ $:$ $\Lambda^{1}\otimes\Lambda^{3}\rightarrow\Lambda^{2}\otimes\Lambda^{2}$
[…]
@@
${\Lambda^{3}\otimes\Lambda^{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\kappa}\scriptstyle{\cong}$${\Lambda^{2}\otimes\Lambda^{2}}$${\Lambda^{1}\otimes\Lambda^{3}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\tilde{\kappa}}\scriptstyle{\cong}$${\mathrm{image}(\kappa)=\mathrm{image}(\tilde{\kappa}).\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
@@
Under the isomorphism (13), the automorphism $\tau\circ\alpha\in\mathrm{End}(\Lambda^{3}\otimes\Lambda^{1})$ identifies as (1,1)(1,-1). In particular, $\tau\circ\alpha$ is an involution, i.e., $(\tau\circ\alpha)\circ(\tau\circ\alpha)=\mathrm{Id}$.
@@ 4.4. Extending $\alpha$ for $r>3$
    By Proposition 4.10, we have $\alpha_{j;i}^{-1}=\tau_{j;i}\circ\alpha_{j;i}\circ\tau_{j:i}$, where $\tau_{j;i}$ transposes the $i$th and $j$th tensor factors.
Example 4.14.
  1. (1)
    For $r=2$, one has $\alpha_{1;2}=\alpha$.
  2. (2)
    If $r>2$, then $\alpha_{1:2}$ is $\alpha\otimes\mathrm{Id}_{T^{r-2}}$.
The significance of $\alpha$ is shown by the following
Lemma 4.15.
Consider $X_{1}\in\Lambda^{3}\otimes\Lambda^{1}$ and $X_{2}\in\Lambda^{1}\otimes\Lambda^{3}$. Then $(X_{1},X_{2})\in\Lambda^{3}\otimes\Lambda^{1}\ \oplus\ \Lambda^{1}\otimes\Lambda^{3}$ lies in the kernel of $\overline{\delta}_{2}$ if and only if
$\alpha_{1;2}(X_{1})=\alpha(X_{1})=-X_{2}.$
Proof.
By Lemma 4.7, $(X_{1},X_{2})\in\Lambda^{3}\otimes\Lambda^{1}\ \oplus\ \Lambda^{1}\otimes\Lambda^{3}$ lies in the kernel of $\overline{\delta}_{2}$ if and only if $\kappa(X_{1})+\tilde{\kappa}(X_{2})=0$. By the definition of $\alpha$, this is equivalent to the condition in the statement, using the identification $\alpha_{1;2}=\alpha$ observed in Example 4.14. ∎
@@ 4.5. Identifying $\ker\overline{\delta}_{r}$
Suppose that $r\geq 2$. Motivated by Proposition 3.10, we seek to calculate $\ker\overline{\delta}_{r}$, where $\overline{\delta}_{r}$ is the morphism introduced in Notation 3.8. The latter has the form given in Lemma 3.9:
@@
Given a polynomial functor $F$, we will use the shorthand ‘element’ to refer to an element of $F(V)$ for some $V\in\mathrm{Ob\ }\mathbf{mod}_{\mathbb{Q}}$, and write $x\in F$ for such an element.
Notation 4.16.
@@
The case $r=2$ has already been established by Lemma 4.15; the general case extends this, as explained below.
The situation can be visualized by using the $1$-skeleton of the (r1)(r-1)-simplex, as in Section 3.3. The elements $X_{i}$ label the black vertices (for $1\leq i\leq r$) and the arrows correspond to the components of $\overline{\delta}_{r}$; a white vertex corresponds to a pair $i<j$.
By the previous identifications, the element $(X_{i})_{1\leq i\leq r}$ lies in the kernel of $\overline{\delta}_{r}$ if and only if, for each $i<j$, the respective images of $X_{i}$ and $X_{j}$ in the corresponding white vertex sum to zero.
Explicitly, the pair $(X_{i},X_{j})$ must lie in the kernel of the corresponding map
$(T^{i-1}\otimes\Lambda^{3}\otimes T^{r-i})\ \oplus\ (T^{j-1}\otimes\Lambda^{3}\otimes T^{r-j})\rightarrow T^{i-1}\otimes\Lambda^{2}\otimes T^{j-i-1}\otimes\Lambda^{2}\otimes T^{r-j}.$
The component maps are induced by $\kappa$ and $\tilde{\kappa}$ applied with respect to the $i$th and $j$th tensor factors. For example, in the case $i=1$, $j=2$, the components identify respectively as $\kappa\otimes\mathrm{Id}_{T^{r-2}}$ and $\tilde{\kappa}\otimes\mathrm{Id}_{T^{r-2}}$.
As in the proof of the case $r=2$ in Lemma 4.15, the previous condition is equivalent to $\alpha_{i;j}(X_{i})=-X_{j}$, by the definition of $\alpha_{i;j}$. ∎
@@
    the underlying inclusion $\mathbf{S}_{(2,1^{2})}\hookrightarrow\Lambda^{3}\otimes\Lambda^{1}\oplus\Lambda^{1}\otimes\Lambda^{3}$ corresponds to the commutative diagram:
@@
By the discussion preceding the statement and using the case $r=2$ of Proposition 4.17, we have an isomorphism
(21) $\Lambda^{3}\otimes\Lambda^{1}\stackrel{{\scriptstyle\cong}}{{\rightarrow}}\ker\overline{\delta}_{2}\subset\Lambda^{3}\otimes\Lambda^{1}\ \oplus\ \Lambda^{1}\otimes\Lambda^{3},$
where the composite is given by $X\mapsto(X,-\alpha X)$.
The action of $\mathfrak{S}_{2}$ on $\ker\overline{\delta}_{2}$ is induced by the transposition of tensor factors $\Lambda^{3}\otimes\Lambda^{1}\stackrel{{\scriptstyle\tau}}{{\rightarrow}}\Lambda^{1}\otimes\Lambda^{3}$. Hence, under the isomorphism (21), this corresponds to the automorphism $X\mapsto-(\tau\circ\alpha)X$ of $\Lambda^{3}\otimes\Lambda^{1}$. By Proposition 4.10, this acts via the diagonal matrix (1,1)(-1,1) (using the identification of Lemma 4.3, as in Proposition 4.10). This proves the first statement and also identifies the inclusion of $\mathbf{S}_{(2,1^{2})}$.
The identification of the embedding of $\Lambda^{4}$ then follows from the graded cocommutativity of the coproduct of $\Lambda^{*}$. ∎
@@ 4.6. A lower bound for $\ker\overline{\delta}_{r}$
We first observe that the terms $\Lambda^{r+2}\boxtimes\mathrm{sgn}_{r}$ and $\mathbf{S}_{(r,1^{2})}\boxtimes\mathrm{triv}_{r}$ are simple and non-isomorphic, hence it suffices to exhibit the appropriate inclusions for the individual factors.
For $\lambda\in\{(1^{r+2}),(r,1^{2})\}$, we have the following useful identification
$\mathrm{Nat}(\mathbf{S}_{\lambda},T^{i-1}\otimes\Lambda^{3}\otimes T^{r-i})=\mathbb{Q}$
for each $1\leq i\leq r$, that is seen as follows (noting that it suffices to show this for $i=1$). Observe that the Young diagram for $(1^{r+2})$ is obtained from that for $(1^{3})$ by adding r1r-1 boxes to the first (and unique) column; likewise, the Young diagram for $(r,1^{2})$ is obtained from that for $(1^{3})$ by adding r1r-1 boxes to the first row. The assertion follows by using Pieri’s rule in conjunction with this observation, together with Proposition 2.4.
Let us fix inclusions (unique up to non-zero scalar multiple, by the above)
$\varphi_{1}$ $:$ $\Lambda^{r+2}\hookrightarrow\Lambda^{3}\otimes T^{r-1}$
$\psi_{1}$ $:$ $\mathbf{S}_{(r,1^{2})}\hookrightarrow\Lambda^{3}\otimes T^{r-1},$
where the subscripts indicate that these correspond to $i=1$. To be concrete, we take $\varphi_{1}$ to be the iterated coproduct for the exterior algebra $\Lambda^{*}$. For $\psi_{1}$, we first choose a (unique up to non-zero scalar multiple) inclusion $\mathbf{S}_{(r,1^{2})}\hookrightarrow\Lambda^{3}\otimes\Gamma^{r-1}$ (see Example 2.10), and then compose with the inclusion $\Lambda^{3}\otimes\Gamma^{r-1}\hookrightarrow\Lambda^{3}\otimes T^{r-1}$ induced by the canonical inclusion $\Gamma^{r-1}\subset T^{r-1}$.
Recall that the group $\mathfrak{S}_{r}$ acts on the domain $(\Lambda^{3}\otimes T^{r-1})\uparrow_{\mathfrak{S}_{r-1}}^{\mathfrak{S}_{r}}$ of $\overline{\delta}_{r}$. In particular, for any $\sigma\in\mathfrak{S}_{r}$, we have the restriction of the action of $\sigma$ to the direct summand $\Lambda^{3}\otimes T^{r-1}$:
$\Lambda^{3}\otimes T^{r-1}\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}T^{i-1}\otimes\Lambda^{3}\otimes T^{r-i}\subset(\Lambda^{3}\otimes T^{r-1})\uparrow_{\mathfrak{S}_{r-1}}^{\mathfrak{S}_{r}}$
where $i=\sigma^{-1}(1)$. (Note the (slightly abusive) usage of the notation $\sigma$.)
We claim that, for given $i$ and $\sigma\in\mathfrak{S}_{r}$ such that $i=\sigma^{-1}(1)$, the following composites:
(22) $\Lambda^{r+2}\stackrel{{\scriptstyle\varphi_{1}}}{{\longrightarrow}}\Lambda^{3}\otimes T^{r-1}\stackrel{{\scriptstyle\mathrm{sgn}(\sigma)\sigma}}{{\longrightarrow}}T^{i-1}\otimes\Lambda^{3}\otimes T^{r-i}$
(23) $\mathbf{S}_{(r,1^{2})}\stackrel{{\scriptstyle\psi_{1}}}{{\longrightarrow}}\Lambda^{3}\otimes T^{r-1}\stackrel{{\scriptstyle\sigma}}{{\longrightarrow}}T^{i-1}\otimes\Lambda^{3}\otimes T^{r-i}$
do not depend on the choice of $\sigma$. This follows easily from the explicit descriptions of $\varphi_{1}$ and $\psi_{1}$ given above.
Hence, for any such choice of $\sigma$, we may define $\varphi_{i}$ to be the composite (22) and $\psi_{i}$ to be the composite (23). (This definition is imposed by the isomorphism that we are trying to establish, due to the terms $\mathrm{sgn}_{r}$ and $\mathrm{triv}_{r}$ appearing in the statement.) By construction, these yield an inclusion
$\Lambda^{r}\oplus\mathbf{S}_{(r,1^{2})}\hookrightarrow(\Lambda^{3}\otimes T^{r-1})\uparrow_{\mathfrak{S}_{r-1}}^{\mathfrak{S}_{r}}.$
To conclude, we must prove that they map to $\ker\overline{\delta}_{r}$, since the construction ensures that the $\mathfrak{S}_{r}$-action is correct (as remarked above).
Hence, by Proposition 4.17, it suffices to prove that, for any $i<j$, the following two diagrams anti-commute:
[…]
and
${\mathbf{S}_{(r,1^{2})}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi_{i}}$$\scriptstyle{\psi_{j}}\scriptstyle{-}$${T^{i-1}\otimes\Lambda^{3}\otimes T^{r-i}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha_{i;j}}$${T^{j-1}\otimes\Lambda^{3}\otimes T^{r-j}.}$
The case $i=1$ and $j=2$ is a straightforward extension of Theorem 4.20. (For the case $\Lambda^{r+2}$, one uses the coassociativity of the coproduct of $\Lambda^{*}$ (denoted $\Delta$ below) to give the factorization of $\varphi_{1}$ as the composition:
$\Lambda^{r+2}\stackrel{{\scriptstyle\Delta}}{{\longrightarrow}}\Lambda^{4}\otimes\Lambda^{r-2}\stackrel{{\scriptstyle\Delta\otimes\Delta}}{{\longrightarrow}}\Lambda^{3}\otimes\Lambda^{1}\otimes(\Lambda^{1})^{\otimes r-2}\cong\Lambda^{3}\otimes T^{r-1}.$
This makes the relationship with the case $r=2$ transparent.)
Finally, for the general case $i<j$, we exploit the relationship between the different $\alpha_{i;j}$ in terms of the $\mathfrak{S}_{r}$-action to reduce to the case $i=1$, $j=2$. This is based on the following commutative diagram, using restrictions of the (left) action of $\mathfrak{S}_{r}$ on $(\Lambda^{3}\otimes T^{r-1})\uparrow_{\mathfrak{S}_{r-1}}^{\mathfrak{S}_{r}}$, as above:
${\Lambda^{3}\otimes T^{r-1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(12)}$$\scriptstyle{(1i)}$${T^{1}\otimes\Lambda^{3}\otimes T^{r-2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(1i)(12)(j2)}$${T^{i-1}\otimes\Lambda^{3}\otimes T^{r-i}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(ij)}$${T^{j-1}\otimes\Lambda^{3}\otimes T^{j-i}}$
(the commutativity is checked directly in $\mathfrak{S}_{r}$).
One has an analogous commutative diagram in which the horizontal maps are replaced by $\alpha_{1;2}$ and $\alpha_{i;j}$ respectively:
${\Lambda^{3}\otimes T^{r-1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha_{1;2}}$$\scriptstyle{(1i)}$${T^{1}\otimes\Lambda^{3}\otimes T^{r-2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(1i)(12)(j2)}$${T^{i-1}\otimes\Lambda^{3}\otimes T^{r-i}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha_{i;j}}$${T^{j-1}\otimes\Lambda^{3}\otimes T^{j-i}.}$
Composing the respective anti-commutative triangles for the case $i=1$, $j=2$ with this commutative square yields the general case $i<j$, by the definition of $\varphi_{i}$ (respectively $\psi_{i}$) in terms of $\varphi_{1}$ (resp. $\psi_{1}$). In the case $\Lambda^{r+2}$, this relies upon the fact that both the vertical maps are induced by permutations of sign 1-1. ∎